# The birth mass function of neutron stars

### v1

@@ A Mass measurements of neutron stars
In Extended Data Tables 1 and 2, we list a total of 90 neutron star mass measurements compiled from the literature. These include 13 double neutron star systems, 27 neutron star-white dwarf binaries, 3 neutron star-main sequence star binaries, 12 “spider” binary pulsar systems (6 redback and 6 black widow binaries), 1 isolated neutron star, 14 X-ray binaries (11 high-mass and 3 low-mass X-ray binaries), 5 gravitational-wave events of compact binary mergers (2 binary neutron star and 3 neutron star-black hole systems).
@@
A key aspect of this work is that we divide neutron stars into two subclasses: recycled and nonrecycled (slow) neutron stars. For most neutron stars, the classification is based on measured spin properties; recycled neutron stars are those with $P\lesssim 100\,\mathrm{ms}$ and $\dot{P}\lesssim 10^{-17}$, whereas slow neutron stars have spin periods of order seconds or longer, and for which no significant accretion-induced spin-up has yet occurred. Moreover, we make the following general assumptions. First, for double neutron star systems and binary neutron star mergers, we assume each consists of a recycled ($m_{r}$) and a slow ($m_{s}$) neutron star, based on the standard formation theory as represented by the double pulsar system PSR J0737--3039A/B[2]. Second, neutron stars in high-mass X-ray binaries are considered to be slow neutron stars; these are neutron stars in the early stage of the recycling process and our assumption is effectively that the amount of accreted mass is negligible. Third, we assume that neutron stars in neutron star-black hole mergers are nonrecycled; this is believed to be the prevailing scenario where the neutron star is the secondly-formed object in the standard isolated binary formation channel[3, 4].
@@
Our division results in 56 recycled neutron stars and 34 slow neutron stars. For the two recycled neutron stars in GW170817 and GW190425, and the unseen companion of PSR J1906+0746 (assumed to be a recycled neutron star), neither $P$ nor $\dot{P}$ are measured. Therefore, we do not include these three $m_{r}$ measurements while accounting for the accreted mass in inferring their birth masses. PSRs J0514--4002A, B2127+11C and J1807--2500B are located in globular clusters, which means that they might have originated from a different dynamical formation channel[5, 6, 7, 8]. Nevertheless we include their companions in the category of slow neutron stars (which is a possibility even for the non-standard dynamical formation channel). Third, it was argued that the companion of PSR J0453+1559 could be a white dwarf born in a thermonuclear electron-capture supernova[9]. We do not expect our analysis results to be significantly affected by the inclusion of these four $m_{s}$ measurements (i.e., companions for three globular-cluster pulsars and PSR J0453+1559).
@@
The input for our hierarchical Bayesian analysis (described in Section C) are individual mass posteriors. For mass measurements reported with symmetric errors in original publications, they are approximated by Gaussian functions; for gravitational-wave events, we use posterior samples that are publicly available as part of relevant publications[10, 11, 12, 13, 14].
@@
Extended Data Table 2: The same as Extended Data Table 1, but for neutron stars in other types of systems. Entries filled with a dash line are inapplicable, while leaving blank indicates an unknown parameter.
@@ B Modelling the accreted masses of recycled pulsars
In this section, we describe our methodology for modelling the accreted mass of a recycled pulsar. Our starting point is the measured properties of recycled pulsars, including the spin period, spin-down rate and mass. The goal is to estimate the accreted mass gained in the recycling process and hence estimate the birth masses for each of these pulsars. Our method involves an analytical approach described in the first part of Section B.1. To fully account for the incomplete understanding of the recycling process, we also consider a phenomenological model as detailed in the second part of Section B.2. As mentioned in Section A, there are 53 recycled neutron stars for which we wish to estimate the birth masses. They are placed in two groups: group $a$ is composed of all recycled pulsars listed in Extended Data Table 1, as well as the isolated millisecond pulsar J0030+0451, and PSR J1903+0327; group $b$ consists of 12 “spider” pulsars, and 2 recycled neutron stars in low-mass X-ray binaries. For pulsars in group $a$, it is generally thought that they might have experienced significant spin-down since the end of the recycling process, whereas pulsars in group $b$ are likely to have completed the accretion not long before [^1] The transitional millisecond pulsar J1023+0038 is an excellent example; its optical variability indicated that it had an accretion disk in 2001[64, 65]..
@@
In the standard binary-recycling paradigm, recycled pulsars start their post-recycling evolution on a spin-up line, defined as $\dot{P}\sim P^{4/3}$, in the $P$-$\dot{P}$ diagram[66, 27]. The exact location of such a spin-up line depends on pulsar-specific properties such as mass, moment of inertia and magnetic inclination angle. For a population of millisecond pulsars, strong observational evidence was recently found for a spin-up line[67]; all pulsars are located to the lower right of this line in the $P$-$\dot{P}$ diagram. For pulsars in group $a$, we account for the pulsar spin-down evolution by speculating where their spin-up lines might lie in the $P$-$\dot{P}$ diagram. For group $b$, it is assumed that the spin-down since the end of the recycling process is negligible, and therefore the accreted mass is linearly correlated with the observed spin period.
@@
The pulsar spin-down evolution can be described in a generic model $\dot{P}\propto P^{2-n}$, where $n$ is the braking index. If magnetic dipole radiation is solely responsible for pulsar spin down, then $n=3$. Observationally, the braking index can be determined via $n=2-(P\ddot{P})/\dot{P}^{2}$ with $\ddot{P}$ being the measured second derivative of the spin period. Currently, the braking index is only measured for a handful of normal pulsars[68], with measured values ranging from 0.9 to 3.15. One theory for $n$ deviating from 3 is the magnetic field decay, which might occur for isolated young (normal) pulsars[69]. Since magnetic field decay is unlikely for (old) recycled pulsars[70], we adopt the standard [^2] We verified that our results are not sensitive to the adopted value of $n$. For example, results with $n=2$ are essentially the same. value of $n=3$.
@@
For a magnetized neutron star accreting matter from a companion star, it may reach an equilibrium spin period, $P_{\rm eq}$, that is given by the Keplerian orbital period at the inner edge of the accretion disk:
$P_{\rm eq}=2\pi\sqrt{\frac{r_{\rm{mag}}^{3}}{Gm}}\frac{1}{\omega_{\rm c}}\,.$ (1)
@@
Here $G$ is the gravitational constant, $m$ is pulsar mass, and $0.25<\omega_{\rm c}\leq 1$ is the critical fastness parameter[66]. The inner radius of the accretion disk $r_{\rm{mag}}=\phi\,r_{\rm{A}}$, with $r_{\rm{A}}$ being the Alfvén radius, which is a characteristic radius at which magnetic stresses dominate the matter flow in the accretion disk:
$r_{\rm{A}}\simeq\left(\frac{B^{2}R^{6}}{\dot{M}\sqrt{2Gm}}\right)^{2/7}\,,$ (2)
@@
where $B$ is the magnetic field strength, $R$ is the pulsar radius, $\dot{M}$ is the accretion rate. The parameter $\phi$ is generally assumed to be in the range between 0.5 and 1.4, depending on the poorly known disk-magnetosphere interactions[66].
@@
where $c$ is the speed of light, $I$ is the pulsar moment of inertia, $\alpha$ is the misalignment angle between the magnetic dipole moment and spin axis. Note that Equation (3) is a convenient approximation to numerical results found in Ref.[71] which takes into account the plasma contribution in the spin-down torque. Inspecting Equations (1-3), one finds that $\dot{P}=AP^{4/3}$, with the coefficient $A$ given by[27]
@@
As can be seen from Equation (4), the spin-up line is not uniquely defined; instead the coefficient $A$ depends on pulsar-specific parameters. In this work, we adopt $A=A_{\rm max}=3.76\times 10^{-15}\,\mathrm{s}^{-4/3}$ as the upper limit for the population of recycled pulsars[67]. We also adopt the following empirical relation for the moment of inertia[72]
$I\simeq 0.237mR^{2}\Bigg{[}1+4.2\frac{m}{\rm{M}_{\odot}}\frac{\rm km}{R}+90\Bigg{(}\frac{m}{\rm{M}_{\odot}}\frac{\rm km}{R}\Bigg{)}^{4}\Bigg{]}\,.$ (5)
The intersection between the spin-down line $\dot{P}\sim P^{2-n}$ and the limiting spin-up line (with $A=A_{\rm max}$) gives a minimum initial spin period for recycled pulsars. However, for some pulsars this minimum period (determined from the intersection of two lines) might be unphysically small, in which case we set the minimum spin period to be $1\,\mathrm{ms}$. This is close to the break-up spin period of neutron stars[73]; also note that in a recent numerical simulation study, the spin period of a neutron star at the end of the recycling process is found to be $\gtrsim 1\,\mathrm{ms}$ for a range of binary evolution calculations[28]. The time spent for a pulsar to evolve from an initial spin period ($P_{0}$) to the current location in the $P$-$\dot{P}$ diagram is
@@
Here if $P_{0}\ll P$ and $n=3$, it reduces to the expression for the characteristic age ($\tau_{\rm c}$). We denote the time for a pulsar to evolve from either the spin-up line or $P_{0}=1\,\mathrm{ms}$ to its current period as the maximum age $\tau_{\rm{max}}$. If $\tau_{\rm{max}}$ is greater than the age of Milky Way $11\,\mathrm{Gyr}$[74, 75], we set it to be $11\,\mathrm{Gyr}$.
In our calculations, we first obtain pulsar intrinsic spin-down rate $\dot{P}_{\rm{int}}$ by correcting the measured $\dot{P}$ for contribution from the Shklovskii effect[76] and Galactic potential[77, 78] using the package GalDynPsr[79]. For three pulsars in globular clusters, PSRs J1910--5959A, B2127$+$11C and J1807--2500B, the contribution from cluster potential is also included[80].
@@
Extended Data Figure 1 shows the locations of 39 recycled pulsars (of group $a$) in the $P$-$\dot{P}$ diagram. Their observed locations with intrinsic $\dot{P}_{\rm{int}}$ are marked by black circles. Orange stars indicate the minimum spin period $P_{\rm lim}$ and its corresponding initial spin-down rate $\dot{P}_{\rm lim}$. For a braking index of 3, black dotted lines ($\dot{P}\sim P^{-1}$) connect the black circles and orange stars. Note that $P_{\rm lim}$ and $\dot{P}_{\rm lim}$ are determined by three constraints: the limiting spin-up line (green line in Extended Data Figure 1), a minimum spin period of $1\,\mathrm{ms}$ (a vertical blue line), and a maximum age of $11\,\mathrm{Gyr}$. For orange stars that do not fall onto the green or blue lines, their locations are a result of the age limit. Light blue stars in this plot indicate plausible places of recycled pulsars at the end of recycling process; since we assume that the age of a recycled pulsar is uniformly distribution in $[0,\tau_{\rm{max}}]$, blue stars correspond to an age of $\tau_{\rm{max}}/2$.
@@ B.1 An analytical model
In our analytical model, we approximate the equilibrium spin period of a neutron star at the end of the recycling process as the minimum spin period accelerated from rest[30]
@@
where $m_{\rm ini}=m-\Delta m$ is the neutron star mass prior to accretion (i.e., its birth mass), $m$ is the measured neutron star mass, and $\Delta m$ is the accreted mass that is being modelled here.
From Equations (1-5) and (7), it can be seen that, the spin-up line and the accreted mass is uniquely determined for a given set of parameters $\{\dot{M},\alpha,R,\phi,\omega_{\rm c}\}$, if we assume that a recycled pulsar starts its rejuvenated life from the spin-up line and spins down due to magnetic dipole braking. Therefore, we model $\Delta m$ in a parameter-agnostic way. For each recycled pulsar, we set the following priors on its recycling parameters. The accretion rate, when expressed in units of the Eddington accretion rate, follows a log uniform distribution between 0.01 and 3; here the upper end is chosen to allow for possible super-Eddington accretion[81]. The magnetic inclination angle $\alpha$ is distributed isotropically, with $\cos\alpha$ uniform between $-1$ and 1. The dimensionless parameters $\phi$ and $\omega_{\rm c}$ are uniformly distributed in the ranges of $[0.5,1.4]$ and $[0.25,1]$, respectively. The pulsar radius follows a Gaussian distribution with a mean of $11.9\,\mathrm{km}$ and a standard deviation of $0.85\,\mathrm{km}$, based on gravitational-wave observations of the first neutron star merger GW170817[82]. We have verified that the effect of using a mass-radius relation for a specific equation of state that is consistent with current observational constraints, e.g., the AP4 equation of state[83], is on the order of $\lesssim 10\%$.
For each pulsar, we generate a sample in the 6-D parameter space $\{m,\dot{M},\alpha,R,\phi,\omega_{\rm c}\}$, and draw a spin-up line (which must lie to the upper left of the pulsar’s current location in the $P$-$\dot{P}$ diagram). Then, the intersection between this spin-up line and the constant magnetic field strength line $\dot{P}\sim P^{-1}$ determines an equilibrium spin period $P_{\rm eq}$. The accreted mass $\Delta m$ is then computed from Equation (7). Note that the statistical uncertainties of $P$ and $\dot{P}$ are ignored, which is an excellent approximation for all pulsars except one globular-cluster pulsar.
For recycled neutron stars in group $b$, we adopt a simple scaling relation (plotted as red dashed line in Extended Data Figure 2) between the accreted mass $\Delta m$ and the observed spin period: $\Delta m/\mathrm{M}_{\odot}=0.2P_{\rm ms}^{-1}$ where $P_{\rm ms}$ is the spin period in ms; see, e.g., equation (6) in ref.[84]. Such a relation can be obtained from Equation (7) in case of $\Delta m\ll m$.

### v2

@@ A Mass measurements of neutron stars
Extended Data Tables 1 and 2 list measurements of the mass of 90 neutron stars compiled from the literature. These include 13 double neutron star systems, 27 neutron star and white dwarf binaries, three neutron star and main sequence star binaries, 12 “spider” binary pulsar systems (six redback and six black widow binaries), one isolated neutron star, 14 X-ray binaries (11 high-mass and three low-mass X-ray binaries), five gravitational-wave events of compact binary mergers (two binary neutron star and three neutron star and black hole systems).
@@
A key aspect of this work is that we divided neutron stars into two subclasses: recycled and non-recycled (slow) neutron stars. For most neutron stars, the classification is based on measured spin properties; recycled neutron stars have period $P\lesssim 100\,\mathrm{ms}$ and $\dot{P}\lesssim 10^{-17}$, whereas slow neutron stars have spin periods of the order of seconds or longer. For the latter, no notable accretion-induced spin-up has yet occurred. Moreover, we made the following general assumptions. First, for double neutron star systems and binary neutron star mergers, we assumed that each consists of a recycled and a slow neutron star, with masses $m_{r}$ and $m_{s}$, based on the standard formation theory as represented by the double pulsar system PSR J0737--3039A/B[2]. Second, neutron stars in high-mass X-ray binaries are considered to be slow neutron stars; these are neutron stars in the early stage of the recycling process and our assumption is effectively that the amount of accreted mass is negligible. Third, we assumed that neutron stars in neutron star and black hole mergers are non-recycled. This is believed to be the prevailing scenario in which the neutron star is the second object formed in the standard isolated binary formation channel[3, 4].
@@
Our division results in 56 recycled neutron stars and 34 slow neutron stars. For the two recycled neutron stars in GW170817 and GW190425, and the unseen companion of PSR J1906+0746 (assumed to be a recycled neutron star), neither $P$ nor $\dot{P}$ have been measured. Therefore, we did not include these three $m_{r}$ measurements when accounting for the accreted mass in inferring their birth masses. PSRs J0514--4002A, B2127+11C and J1807--2500B are in globular clusters, which means that they may have originated from a different dynamical formation channel[5, 6, 7, 8]. Nevertheless, we included their companions in the category of slow neutron stars (which is a possibility even for the non-standard dynamical formation channel). Third, it was argued that the companion of PSR J0453+1559 could be a white dwarf born in a thermonuclear electron-capture supernova[9]. We do not expect our analysis results to be substantially affected by the inclusion of these four $m_{s}$ measurements (companions for three globular-cluster pulsars and PSR J0453+1559).
@@
The input for our hierarchical Bayesian analysis (described in Section C in Methods) are individual mass posteriors. For mass measurements reported with symmetric errors in original publications, they are approximated by Gaussian functions; for gravitational-wave events, we used posterior samples that are publicly available in the relevant publications[10, 11, 12, 13, 14].
@@
[…]
@@ B Modelling the accreted masses of recycled pulsars
In this section, we describe our methodology for modelling the accreted mass of a recycled pulsar. Our starting point is the measured properties of recycled pulsars, including the spin period, spin-down rate and mass. The goal is to estimate the accreted mass gained in the recycling process and hence estimate the birth masses for each of these pulsars. Our method involves an analytical approach described in the first part of Section B.1. To fully account for the incomplete understanding of the recycling process, we also consider a phenomenological model as detailed in the second part of Section B.2. As mentioned in Section A, we aimed to estimate the birth masses of 53 recycled neutron stars. They are placed in two groups: group $a$ was composed of all recycled pulsars listed in Extended Data Table 1 as well as the isolated millisecond pulsar J0030+0451 and PSR J1903+0327; group $b$ consists of 12 “spider” pulsars, and two recycled neutron stars in low-mass X-ray binaries. It is generally thought that the pulsars in group $a$ may have experienced notable spin-down since the end of the recycling process, whereas the pulsars in group $b$ probably completed accretion not long before111 The transitional millisecond pulsar J1023+0038 is an excellent example. Its optical variability indicates that it had an accretion disk in 2001[64, 65]..
@@
In the standard binary-recycling paradigm, recycled pulsars start their post-recycling evolution on a spin-up line, defined as $\dot{P}\propto P^{4/3}$, in the $P$ versus $\dot{P}$ diagram[66, 27]. The exact location of such a spin-up line depends on pulsar-specific properties such as mass, moment of inertia and magnetic inclination angle. For a population of millisecond pulsars, strong observational evidence was recently found for a spin-up line[67]. All pulsars were located to the lower right of this line in the $P$-$\dot{P}$ diagram. For pulsars in group $a$, we accounted for the pulsar spin-down evolution by speculating where their spin-up lines might lie in the $P$-$\dot{P}$ diagram. For group $b$, we assumed that the spin-down since the end of the recycling process is negligible and, therefore, that the accreted mass is linearly correlated with the observed spin period.
@@
The pulsar spin-down evolution can be described in a generic model $\dot{P}\propto P^{2-n}$, where $n$ is the braking index. If magnetic dipole radiation is solely responsible for pulsar spin down, then $n=3$. Observationally, the braking index can be determined from $n=2-(P\ddot{P})/\dot{P}^{2}$ with $\ddot{P}$ being the measured second derivative of the spin period. At present, the braking index has been measured only for a handful of normal pulsars[68], with measured values ranging from 0.9 to 3.15. One theory for $n$ deviating from 3 is the magnetic field decay, which could occur for isolated young (normal) pulsars[69]. As magnetic field decay is unlikely for (old) recycled pulsars[70], we adopted the standard [^2] We verified that our results are not sensitive to the adopted value of $n$. For example, results with $n=2$ are essentially the same. value of $n=3$.
@@
A magnetized neutron star accreting matter from a companion star may reach an equilibrium spin period $P_{\rm eq}$ that is given by the Keplerian orbital period at the inner edge of the accretion disk:
$P_{\rm eq}=2\pi\sqrt{\frac{r_{\rm{mag}}^{3}}{Gm}}\frac{1}{\omega_{\rm c}}\,,$ (1)
@@
were $G$ is the gravitational constant, $m$ is pulsar mass, and $0.25<\omega_{\rm c}\leq 1$ is the critical fastness parameter[66]. The inner radius of the accretion disk $r_{\rm{mag}}=\phi\,r_{\rm{A}}$, with $r_{\rm{A}}$ being the Alfvén radius, which is a characteristic radius at which magnetic stresses dominate the matter flow in the accretion disk:
$r_{\rm{A}}\approx\left(\frac{B^{2}R^{6}}{\dot{M}\sqrt{2Gm}}\right)^{2/7}\,,$ (2)
@@
where $B$ is the magnetic field strength, $R$ is the pulsar radius and $\dot{M}$ is the accretion rate. The parameter $\phi$ is generally assumed to be in the range between 0.5 and 1.4, depending on the poorly known disk-magnetosphere interactions[66].
@@
where $c$ is the speed of light, $I$ is the pulsar moment of inertia and $\alpha$ is the misalignment angle between the magnetic dipole moment and the spin axis. Note that Equation (3) is a convenient approximation of numerical results in ref.[71], which takes into account the plasma contribution in the spin-down torque. Inspecting Equations (1-3), one finds that $\dot{P}=AP^{4/3}$, with the coefficient $A$ given by[27]
@@
As can be seen from Equation (4), the spin-up line is not uniquely defined. Instead, the coefficient $A$ depends on pulsar-specific parameters. In this work, we adopt $A=A_{\rm max}=3.76\times 10^{-15}\,\mathrm{s}^{-4/3}$ as the upper limit for the population of recycled pulsars[67]. We also adopted the following empirical relation for the moment of inertia[72]
$I\approx 0.237mR^{2}\Bigg{[}1+4.2\frac{m}{\rm{M}_{\odot}}\frac{\rm km}{R}+90\Bigg{(}\frac{m}{\rm{M}_{\odot}}\frac{\rm[km]}{R}\Bigg{)}^{4}\Bigg{]}\,.$ (5)
The intersection between the spin-down line $\dot{P}\propto P^{2-n}$ and the limiting spin-up line (with $A=A_{\rm max}$) gives a minimum initial spin period for recycled pulsars. However, for some pulsars this minimum period (determined from the intersection of two lines) was unphysically small, in which case we set the minimum spin period to $1\,\mathrm{ms}$. This is close to the break-up spin period of neutron stars[73]. Also note that in a recent numerical simulation study, the spin period of a neutron star at the end of the recycling process was found to be $\gtrsim 1\,\mathrm{ms}$ for a range of binary evolution calculations[28]. The time spent for a pulsar to evolve from an initial spin period ($P_{0}$) to its current location in the $P$-$\dot{P}$ diagram is
@@
Here if $P_{0}\ll P$ and $n=3$, this reduces to the expression for the characteristic age ($\tau_{\rm c}$). We denote the time for a pulsar to evolve from either the spin-up line or $P_{0}=1\,\mathrm{ms}$ to its current period as the maximum age $\tau_{\rm{max}}$. If $\tau_{\rm{max}}$ is greater than the age of Milky Way, $11\,\mathrm{Gyr}$ (refs.[74, 75]), we set it to $11\,\mathrm{Gyr}$.
In our calculations, we first obtained pulsar intrinsic spin-down rate $\dot{P}_{\rm{int}}$ by correcting the measured $\dot{P}$ with contribution from the Shklovskii effect[76] and Galactic potential[77, 78] using the package GalDynPsr[79]. For three pulsars in globular clusters, PSRs J1910--5959A, B2127$+$11C and J1807--2500B, the contribution from cluster potential was also included[80].
@@
Extended Data Figure 1 shows the locations of 39 recycled pulsars (in group $a$) in the $P$-$\dot{P}$ diagram. Their observed locations with intrinsic $\dot{P}_{\rm{int}}$ are marked by black circles. Orange stars indicate the minimum spin period $P_{\rm lim}$ and its corresponding initial spin-down rate $\dot{P}_{\rm lim}$. For a braking index of 3, black dotted lines ($\dot{P}\propto P^{-1}$) connect the black circles and orange stars. Note that $P_{\rm lim}$ and $\dot{P}_{\rm lim}$ are determined by three constraints: the limiting spin-up line (green line in Extended Data Figure 1), a minimum spin period of $1\,\mathrm{ms}$ (a vertical blue line), and a maximum age of $11\,\mathrm{Gyr}$. For orange stars that do not fall onto the green or blue lines, their locations are a result of the age limit. Light blue stars in this plot indicate plausible places of recycled pulsars at the end of recycling process. As we assumed that the age of a recycled pulsar is uniformly distributed in $[0,\tau_{\rm{max}}]$, blue stars correspond to an age of $\tau_{\rm{max}}/2$.
@@ B.1 An analytical model
In our analytical model, we approximate the equilibrium spin period of a neutron star at the end of the recycling process as the minimum spin period accelerated from rest[30]:
@@
where $m_{\rm ini}=m-\Delta m$ is the neutron-star mass before accretion (its birth mass), $m$ is the measured neutron-star mass and $\Delta m$ is the accreted mass that is being modelled here.
From Equations (1)-(5) and (7), it can be seen that, the spin-up line and the accreted mass is uniquely determined for a given set of parameters $\{\dot{M},\alpha,R,\phi,\omega_{\rm c}\}$, if we assume that a recycled pulsar starts its rejuvenated life from the spin-up line and spins down due to magnetic dipole braking. Therefore, we modelled $\Delta m$ in a parameter-agnostic way. For each recycled pulsar, we set the following priors on its recycling parameters. The accretion rate, when expressed in units of the Eddington accretion rate, follows a log-uniform distribution between 0.01 and 3. The upper end was chosen to allow for possible super-Eddington accretion[81]. The magnetic inclination angle $\alpha$ is distributed isotropically, with $\cos\alpha$ uniform between $-1$ and 1. The dimensionless parameters $\phi$ and $\omega_{\rm c}$ are uniformly distributed in the ranges $[0.5,1.4]$ and $[0.25,1]$, respectively. The pulsar radius follows a Gaussian distribution with a mean of $11.9\,\mathrm{km}$ and a standard deviation of $0.85\,\mathrm{km}$, based on gravitational-wave observations of the first neutron star merger GW170817[82]. We have verified that the effect of using a mass-radius relation for a specific equation of state that is consistent with current observational constraints, for example, the AP4 equation of state[83], is of the order of $\lesssim 10\%$.
For each pulsar, we generated a sample in the six-dimensional parameter space $\{m,\dot{M},\alpha,\newline R,\phi,\omega_{\rm c}\}$, and draw a spin-up line (which must lie to the upper left of the current location of pulsar in the $P$-$\dot{P}$ diagram). Then, the intersection between this spin-up line and the constant magnetic field strength line $\dot{P}\propto P^{-1}$ determines an equilibrium spin period $P_{\rm eq}$. The accreted mass $\Delta m$ was then computed from Equation (7). Note that the statistical uncertainties of $P$ and $\dot{P}$ were ignored, which is an excellent approximation for all pulsars except one globular-cluster pulsar.
For recycled neutron stars in group $b$, we adopted a simple scaling relation (plotted as the red dashed line in Extended Data Figure 2) between the accreted mass $\Delta m$ and the observed spin period: $\Delta m/\mathrm{M}_{\odot}=0.2P_{\rm ms}^{-1}$ where $P_{\rm ms}$ is the spin period in milliseconds; see, for example, equation (6) in ref.[84]. Such a relation can be obtained from Equation (7) when $\Delta m\ll m$.